MMS Corp borrows $1,650,000 today for a new building. The loan is an equal principal payment loan with an APR of 6.5% compounded monthly. Payments are due monthly and the term of the loan is 9 years.

a) The interest component of the payment due in month 16 is?

b) The payment due in month 16 is?

c)The current portion of debt in month 16 is?
Answer #1

a)

A B C D E F Year Payments Interest Principal Ending Bal. OTT $1,650,000.00 1 $20,220 $8,937.50 $11,282 $1,638,717.55 21 $20,2

The interest component of the payment due in month 16 is $7,985.

b)

Rate = 6.5%/12 = 0.54%
Period = 9 * 12 = 108
Principal = $1,650,000
Monthly payment =
PMT(0.54%,108,-1650000)
= $20,200

c)
Amount of loan = $1,650,000
Number of payments = 9 * 12 = 108
Equal installment amount = $1,650,000 / 108 = $15,277.77778
Current portion of debt in the month 16 = $15,277.77778 * 12
= $183,333.33 or $183,333.